# references.bib

### v1

@@ 4. Fully discrete formulation
In this section, we integrate the semi-discrete system 8 in time to develop a fully discrete formulation of the thermal shallow water equations. Temporal conservation is also analysed.
@@ 4.1. Poisson integrator
The semi-discrete system 8 is a Poisson system of the form
@@
where the discrete variational derivatives of the Hamiltonian are exactly integrated in time [BauerCotter2018, Eldred2019]. By construction of the Poisson integrator [Cohen2011], applying 23 to the semi-discrete system 8 yields temporal energy conservation. Unfortunately the chosen Poisson integrator [Cohen2011] does not generalise to cubic Casimirs, such as discrete entropy, and we are not aware of any time integrator that preserves higher order Casimirs in general for non-canonical Hamiltonian systems.
@@ 4.2. Temporal discretisation
Without loss of generality, let $J^{n}\in(0,1]$. We find temporal nodal values of $b_{h}^{n}\in\mathbb{V}_{2}$ using $\varphi_{h}^{n},B_{h}^{n}\in\mathbb{V}_{2}$ in the semi-discrete relation 10d. Then, $b_{h}^{n-1/2}=(b_{h}^{n-1}+b_{h}^{n})/2$ is the mean value of a linear polynomial in time for $t\in J^{n}$.
@@
and $[b_{h}b_{h}]^{n-1/2}=(b_{h}^{n-1}b_{h}^{n-1}+b_{h}^{n}b_{h}^{n})/2$ is the mean value of the square of $b_{h}$. Quadratures up to the degree required are used to evaluate $\boldsymbol{F}^{n}_{h}\in\mathbb{V}_{1}$, $\Phi^{n}_{h},\vartheta^{n}_{h}\in\mathbb{V}_{2}$, $q_{h}^{n-1/2}\in\mathbb{V}_{0}$ exactly in time using $\boldsymbol{u}_{h}^{n},\varphi_{h}^{n},B_{h}^{n}$ in 10 [Eldred2019, BauerCotter2018].
@@ 4.3. Temporal conservation
Discrete energy is conserved point-wise in time by construction of the Poisson integrator [Cohen2011]. Taking $\psi_{\boldsymbol{u}_{h}}=\boldsymbol{F}^{n}_{h}$, $\phi_{\varphi_{h}}=\Phi_{h}^{n}$, $\phi_{B_{h}}=\vartheta_{h}^{n}$ in 24 yields the temporal approximation of 14 as
$\int_{J^{n}}\tau\frac{\partial\mathcal{H}_{h}}{\partial t}=\mathcal{H}_{h}^{n}-\mathcal{H}_{h}^{n-1}=\left(\boldsymbol{u}_{h}^{n}-\boldsymbol{u}_{h}^{n-1},\boldsymbol{F}_{h}^{n}\right)+\left(\varphi_{h}^{n}-\varphi_{h}^{n-1},\Phi_{h}^{n}\right)+\left(B_{h}^{n}-B_{h}^{n-1},\vartheta_{h}^{n}\right)=0,$ (26)
which is zero since skew symmetry is maintained. Proving 24b holds point-wise in space and time yields discrete mass conservation.
@@
The continuity equation 24b holds point-wise in space and time.
@@
Assuming continuity of time, 8b shows $(\partial_{t}\varphi_{h}+\nabla\cdot\boldsymbol{F}_{h})(x,\cdot)\in\mathbb{V}_{2}$ holds point-wise in space. Interpolation of the initial condition yields $\varphi_{h}^{0}\in\mathbb{V}_{2}$. Since $\varphi_{h}^{n},\nabla\cdot\boldsymbol{F}_{h}^{n}\in\mathbb{V}_{2}$ for $n=1,2,\dots,N$, it is clear to see $(\varphi_{h}^{n}-\varphi_{h}^{n-1}+\tau\nabla\cdot\boldsymbol{F}_{h}^{n})\in\mathbb{V}_{2}$ holds point-wise in time. ∎
A key novelty of the fully discrete formulation in 24 is that the diagnostic buoyancy is represented as a continuous linear polynomial in time. Consequently, discrete entropy is a continuous cubic polynomial in time, which is not conserved point-wise in time by the chosen Poisson integrator [Cohen2011]. We now prove the loss in exact entropy conservation depends on the accuracy of the temporal approximation.
@@
Subsequent to Proposition 4.2, the relation in 17 is not discretely preserved point-wise in time. To demonstrate the importance of 25 in regards to discrete entropy conservation, we consider the temporal approximation of 21 as
$\int_{J^{n}}\tau\frac{\partial\mathcal{S}_{h}}{\partial t}=\mathcal{S}_{h}^{n}-\mathcal{S}_{h}^{n-1}=\frac{1}{2}\left([b_{h}b_{h}]^{n-1/2},\nabla\cdot\boldsymbol{F}_{h}^{n}\right)-\frac{1}{2}\left(\widetilde{b_{h}^{n}}b_{h}^{n-1/2},\nabla\cdot\boldsymbol{F}_{h}^{n}\right)-s_{\mathrm{up}}(\boldsymbol{F}_{h}^{n},b_{h}^{n-1/2},b_{h}^{n-1/2}).$ (28)
Using $\widetilde{\phi}_{h}=\nabla\cdot\boldsymbol{F}_{h}^{n}$ in 25 yields $\partial\mathcal{S}_{h}/\partial t=0$ when $\alpha=0$, and ${\partial\mathcal{S}_{h}}/{\partial t}<0$ when $\alpha>0$. We emphasise the choice of $\widetilde{b_{h}^{n}}$ means no spatial or temporal entropy conservation errors arise in the forcing terms. Thus, the only source of entropy conservation error is the approximation of temporal derivatives. Numerical experiments in Section 5 below support this analysis.
@@ 4.4. Constrained formulation
where $\lambda\in\mathbb{R}$ is a Lagrange multiplier. Diagnosing $b_{h}^{n}$ using 30 instead of 10d means the analysis of temporal entropy conservation in Section 4.3 is trivial, since we explicitly enforce $\tau\partial\mathcal{S}_{h}/\partial t=\mathcal{S}_{h}^{n}-\mathcal{S}_{h}^{n-1}=0$ point-wise in time. It follows from Proposition 4.2 that $\lambda$ is sufficiently small and corrects an $\mathcal{O}(\tau)$ drift in exact entropy conservation. Further, using 30 perturbs $b_{h}^{n}$ everywhere in 24. Since skew symmetry is maintained independent of $b_{h}^{n}$, discrete energy conservation holds as in 26 by construction of the chosen Poisson integrator [Cohen2011].
@@ 4.5. Quasi-Newton approach
The fully discrete system 24 is an implicit non-linear problem, for which using a quasi-Newton approach to approximate the Jacobian can reduce computational expense. The Jacobian is computed once per time step as
@@
where $\omega_{h}^{n-1}\in\mathbb{X}_{01}$ is diagnosed by setting $\omega_{h}^{n-1}=q_{h}^{n-1}\varphi_{h}^{n-1}$ in 10e. Equation 31 originates from an implicit midpoint discretisation of the linear thermal shallow water equations, in which fast linear waves are accounted for and non-linear coupling terms are omitted [Eldred2019, Lee2022, BauerCotter2018, Wimmer2020]. The linearisation in the current study differs from other works [Eldred2019, Lee2022] as we consider perturbations from the solution at $t^{n-1}$ by using $\varphi_{h}^{n-1},b_{h}^{n-1}$ in 31 instead of the initial usual mean flow state. In doing so, we account for turbulent states which significantly deviate from the resting steady state as simulations progress. Including upwinded numerical fluxes may result in solver convergence issues when simulating a mature turbulent state (see Section 5.2). In these situations, the Jacobian is recomputed at every non-linear iteration and includes approximated centred and upwinded fluxes.
@@ 4.6. Implementation remarks
    Evaluate $\boldsymbol{F}^{n}_{h}\in\mathbb{V}_{1}$, $\Phi^{n}_{h},\vartheta^{n}_{h},b_{h}^{n}\in\mathbb{V}_{2}$, $q_{h}^{n-1/2}\in\mathbb{V}_{0}$ exactly in time via 10.
@@
    Construct $b_{h}^{n-1/2}$ and $[b_{h}b_{h}]^{n-1/2}$, and diagnose $\widetilde{b_{h}^{n}}$ via 25.
@@
which converges to $\operatorname{sign}(x)$ as $\epsilon\rightarrow 0$ for $\epsilon\geq 0$. However 32 is a non-differentiable function that may cause issues in numerical solvers [Badia2017]. An alternative approach is to approximate $\operatorname{sign}(x)$ by a smooth function [Badia2017] such as
@@
which also converges to $\operatorname{sign}(x)$ as $\epsilon\rightarrow 0$. We refer to 32 and 33 as the hard and soft signum functions respectively, and explore their effects in the numerical experiments below. The value of $\epsilon$ is based on values of $||\boldsymbol{F}_{h}^{n}\cdot\boldsymbol{n}^{+}||$ from experiments using the centred scheme.
@@ 5.1. Convergence test
Convergence of the centred scheme is tested using the relative $L^{2}$ norm between the initial and final solutions in the absence of stabilisation [Eldred2019]. Fig. 1(a) shows the expected third and fourth order convergence for a range of spatial and temporal resolutions. For each line in Fig. 1, ${h}$ and ${\tau}$ reduce simultaneously at rate $C=\mathrm{CFL}/p^{2}$, where $\mathrm{CFL}=0.2$ for $p=1$ and $\mathrm{CFL}=0.1$ for $p=2$. Similar convergence rates are observed in Fig. 1(b) for the upwinded scheme.
@@ 5.2. Case study 1: thermal instability
We now consider transient case studies to confirm adequate conservation of system invariants under turbulent flow conditions and over long time periods relative to the onset of geostrophic turbulence. Discrete conservation is demonstrated using normalised values of discrete energy 13, entropy 16 and mass ($\mathcal{M}_{h}(t)=\int_{\Omega}\varphi_{h}$) on a logarithmic scale over time. The change in entropy due to forcing terms, $\partial\mathcal{S}_{h}/\partial t$ in 28, is also shown. In all simulations, the tolerance of the non-linear solver is $10^{-12}$, and conservation errors up to $\mathcal{O}(10^{-12})$ are reasonable.
The first transient case study involves a thermal instability [Eldred2019, Gouzien2017, Zeitlin2018, Kurganov2021]. Other studies simulate this test case using an upwinded finite volume scheme [Kurganov2021, Gouzien2017, Zeitlin2018]. A previous finite element study shows energy conservation at early times [Eldred2019], but entropy conservation, the evolution of complex non-linear flows, and the effect of upwinding is not considered. We extend these works using the novel centred and upwinded fluxes to simulate the thermal instability test case a over long time period while examining conservation of system invariants.
@@
Fig. 2 shows the evolution of the thermal instability for the centred and upwinded schemes. Similar to other studies that also consider this test case [Eldred2019, Gouzien2017], the initial growth evident in Fig. 2(a)–(d) reflects the wave number of 4 used in 36, and non-linear saturation of the instability occurs as the simulation progresses (Fig. 2(e)–(h)). Comparing the buoyancy field at $t=25$ (Fig. 2(a,b)) that obtained at $t=100$ (Fig. 2(g,h)) shows a large variation in buoyancy. Regardless, both the centred and upwinded schemes resolve the highly non-linear flow that develops as the simulation progresses. This is a direct consequence of the new linearised Jacobian 31, which allows for robust convergence for both centred and upwinded fluxes in the presence of significant variations from the initial mean flow state. As expected, including upwinded fluxes (Fig. 2(b,d,f,g)) suppresses spurious oscillations which arise in the absence of stabilisation (Fig. 2(a,c,e,g)).
@@
We now assess the ability of the centred and upwinded schemes to conserve system invariants in the presence of turbulent flow. In the absence of stabilisation, Fig. 3(a,b) shows energy and mass conservation, while Fig. 3(c) demonstrates the entropy conservation error increases to $\mathcal{O}(10^{-8})$ over time even though ${\partial\mathcal{S}_{h}}/{\partial t}$ is machine zero for the entire duration (Fig. 3(d)). This supports the analysis in Section 4.3 above that entropy is not conserved point-wise in time. While an entropy conservation error of $\mathcal{O}(10^{-8})$ may seem large relative to energy and mass, it is comparable with other studies that consider temporal conservation of quadratic tracer invariants [Lee2022]. Thus, we conclude the centred scheme stably resolves turbulent dynamics over long periods of time and conserves system invariants.
@@
For the upwinded scheme, the spikes in energy and mass evident in Fig. 3(a,b) correspond to instances when the non-linear solver fails to converge. Values of $\epsilon=10^{-3}$ and $\epsilon=10^{-4}$ are chosen to approximate $\operatorname{sign}(x)$ in the the upwinded scheme since $||\boldsymbol{F}_{h}^{n}\cdot\boldsymbol{n}^{+}||\approx 0.1$ for the centred scheme. When using the hard signum function with $\epsilon=10^{-4}$, solver convergence issues arise at $t\approx 45$. In contrast, using the soft signum function for $\epsilon=10^{-4}$ and $\epsilon=10^{-3}$ delays the onset of solver convergence issues to $t\approx 70$ and $t\approx 98$ respectively. Therefore, using the soft signum function and increasing $\epsilon$ assists solver convergence, in particular conservation of energy and mass, as turbulent flow dynamics evolve.
@@
Figure 3. Conservation errors for the thermal instability case study. (a,b,c) show normalised values of energy, mass and entropy, and (d) illustrates the change in entropy due to forcing terms 21. Each figure compares the centred and upwinded schemes for different approximations of signum and $\epsilon$ values. The vertical axis is logarithmic and the horizontal axis shows unit-less time. Parameters: $n_{s}=192$ spatial elements, $p=1$ spatial finite elements, $\mathrm{CFL}=0.2$.

### v2

@@ 4. Fully discrete formulation
In this section, we integrate the semi-discrete system 10 in time to develop a fully discrete formulation of the thermal shallow water equations. Temporal conservation is also analysed.
@@ 4.1. Poisson integrator
Since the semi-discrete system 10 originates from a skew symmetric formulation of the thermal shallow water equations, it is a Poisson system of the form
@@
where the discrete variational derivatives of the Hamiltonian are exactly integrated in time [BauerCotter2018, Eldred2019]. By construction of the Poisson integrator [Cohen2011], applying 25 to the semi-discrete system 10 yields temporal energy conservation. Unfortunately the chosen Poisson integrator [Cohen2011] does not generalise to cubic Casimirs, such as discrete entropy, and we are not aware of any time integrator that preserves higher order Casimirs in general for non-canonical Hamiltonian systems.
@@ 4.2. Temporal discretisation
We find temporal nodal values of $b_{h}^{n}\in\mathbb{V}_{2}$ using $\varphi_{h}^{n},B_{h}^{n}\in\mathbb{V}_{2}$ in the semi-discrete relation 12d. Then, $b_{h}^{n-1/2}=(b_{h}^{n-1}+b_{h}^{n})/2$ is the mean value of a linear polynomial in time for $t\in J^{n}$.
@@
and $[b_{h}b_{h}]^{n-1/2}=(b_{h}^{n-1}b_{h}^{n-1}+b_{h}^{n}b_{h}^{n})/2$ is the mean value of the square of $b_{h}$. Quadratures up to the degree required are used to evaluate $\boldsymbol{F}^{n}_{h}\in\mathbb{V}_{1}$, $\Phi^{n}_{h},\vartheta^{n}_{h}\in\mathbb{V}_{2}$, $q_{h}^{n-1/2}\in\mathbb{V}_{0}$ exactly in time using $\boldsymbol{u}_{h}^{n},\varphi_{h}^{n},B_{h}^{n}$ in 12 [Eldred2019, BauerCotter2018].
@@ 4.3. Temporal conservation
Discrete energy is conserved pointwise in time by construction of the Poisson integrator [Cohen2011]. Taking $\psi_{\boldsymbol{u}_{h}}=\boldsymbol{F}^{n}_{h}$, $\phi_{\varphi_{h}}=\Phi_{h}^{n}$, $\phi_{B_{h}}=\vartheta_{h}^{n}$ in 26 yields the temporal approximation of 16 as
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{J^{n}}\frac{\partial\mathcal{H}_{h}}{\partial t}=\mathcal{H}_{h}^{n}-\mathcal{H}_{h}^{n-1}=\left(\boldsymbol{u}_{h}^{n}-\boldsymbol{u}_{h}^{n-1},\boldsymbol{F}_{h}^{n}\right)+\left(\varphi_{h}^{n}-\varphi_{h}^{n-1},\Phi_{h}^{n}\right)+\left(B_{h}^{n}-B_{h}^{n-1},\vartheta_{h}^{n}\right)=0,}$ (28)
which is zero since skew symmetry is maintained. Proving 26b holds pointwise in space and time yields discrete mass conservation.
@@
The continuity equation 26b holds pointwise in space and time.
@@
Assuming continuity of time, 10b shows $(\partial_{t}\varphi_{h}+\nabla\cdot\boldsymbol{F}_{h})(x,\cdot)\in\mathbb{V}_{2}$ holds pointwise in space. Interpolation of the initial condition yields $\varphi_{h}^{0}\in\mathbb{V}_{2}$. Since $\varphi_{h}^{n},\nabla\cdot\boldsymbol{F}_{h}^{n}\in\mathbb{V}_{2}$ for $n=1,2,\dots,N$, it is clear to see $(\varphi_{h}^{n}-\varphi_{h}^{n-1}+\tau\nabla\cdot\boldsymbol{F}_{h}^{n})\in\mathbb{V}_{2}$ holds pointwise in time. ∎
A key novelty of the fully discrete formulation in 26 is that the diagnostic buoyancy is represented as a continuous linear polynomial in time. Consequently, discrete entropy is a continuous cubic polynomial in time, which is not conserved pointwise in time by the chosen Poisson integrator [Cohen2011]. We now prove the loss in exact entropy conservation depends on the accuracy of the temporal approximation.
@@
Subsequent to Proposition 4.2, it should be stated that the relation in 19 is not discretely preserved pointwise in time. To demonstrate the importance of 27 in regards to discrete entropy conservation, we consider the temporal approximation of 23 as
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{J^{n}}\frac{\partial\mathcal{S}_{h}}{\partial t}=\mathcal{S}_{h}^{n}-\mathcal{S}_{h}^{n-1}=\frac{1}{2}\left([b_{h}b_{h}]^{n-1/2},\nabla\cdot\boldsymbol{F}_{h}^{n}\right)-\frac{1}{2}\left(\widetilde{b_{h}^{n}}b_{h}^{n-1/2},\nabla\cdot\boldsymbol{F}_{h}^{n}\right)-s_{\mathrm{up}}(\boldsymbol{F}_{h}^{n},b_{h}^{n-1/2},b_{h}^{n-1/2}).}$ (30)
Using $\widetilde{\phi}_{h}=\nabla\cdot\boldsymbol{F}_{h}^{n}$ in 27 yields $\partial\mathcal{S}_{h}/\partial t=0$ when $\alpha=0$, and ${\partial\mathcal{S}_{h}}/{\partial t}<0$ when $\alpha>0$. We emphasise the choice of $\widetilde{b_{h}^{n}}$ means no spatial or temporal entropy conservation errors arise in the forcing terms. Thus, the only source of entropy conservation error is the approximation of temporal derivatives. Numerical experiments in Section 5 below support this analysis.
@@ 4.4. Constrained formulation
where $\lambda\in\mathbb{R}$ is a Lagrange multiplier. Diagnosing $b_{h}^{n}$ using 32 instead of 12d means the analysis of temporal entropy conservation in Section 4.3 is trivial, since we explicitly enforce $\tau\partial\mathcal{S}_{h}/\partial t=\mathcal{S}_{h}^{n}-\mathcal{S}_{h}^{n-1}=0$ pointwise in time. It follows from Proposition 4.2 that $\lambda$ is sufficiently small and corrects an $\mathcal{O}(\tau)$ drift in exact entropy conservation. Further, using 32 perturbs $b_{h}^{n}$ everywhere in 26. Since skew symmetry is maintained independent of $b_{h}^{n}$, discrete energy conservation holds as in 28 by construction of the chosen Poisson integrator [Cohen2011].
@@ 4.5. Quasi-Newton approach
The fully discrete system 26 is an implicit non-linear problem, for which using a quasi-Newton approach to approximate the Jacobian can reduce computational expense. The Jacobian is computed once per time step as
@@
where $\omega_{h}^{n-1}\in\mathbb{X}_{01}$ is diagnosed by setting $\omega_{h}^{n-1}=q_{h}^{n-1}\varphi_{h}^{n-1}$ in 12e. Equation 33 originates from an implicit midpoint discretisation of the linear thermal shallow water equations, in which fast linear waves are accounted for and non-linear coupling terms are omitted [Eldred2019, Lee2022, BauerCotter2018, Wimmer2020]. The linearisation in the current study differs from other works [Eldred2019, Lee2022] as we consider perturbations from the solution at $t^{n-1}$ by using $\varphi_{h}^{n-1},b_{h}^{n-1}$ in 33 instead of the initial usual mean flow state. In doing so, we account for turbulent states which significantly deviate from the resting steady state as simulations progress. Including upwinded numerical fluxes may result in solver convergence issues when simulating a mature turbulent state (see Section 5.2). In these situations, the Jacobian is recomputed at every non-linear iteration and includes approximated centred and upwinded fluxes.
@@ 4.6. Implementation remarks
    Evaluate $\boldsymbol{F}^{n}_{h}\in\mathbb{V}_{1}$, $\Phi^{n}_{h},\vartheta^{n}_{h},b_{h}^{n}\in\mathbb{V}_{2}$, $q_{h}^{n-1/2}\in\mathbb{V}_{0}$ exactly in time via 12.
@@
    Construct $b_{h}^{n-1/2}$ and $[b_{h}b_{h}]^{n-1/2}$, and diagnose $\widetilde{b_{h}^{n}}$ via 27.
@@
which converges to $\operatorname{sign}(x)$ as $\epsilon\rightarrow 0$ for $\epsilon\geq 0$. However 34 is a non-differentiable function that may cause issues in numerical solvers [Badia2017]. An alternative approach is to approximate $\operatorname{sign}(x)$ by a smooth function [Badia2017] such as
@@
which also converges to $\operatorname{sign}(x)$ as $\epsilon\rightarrow 0$. We refer to 34 and 35 as the hard and soft signum functions respectively, and explore their effects in the numerical experiments below. The value of $\epsilon$ is based on values of $||\boldsymbol{F}_{h}^{n}\cdot\boldsymbol{n}^{+}||$ from experiments using the centred scheme.
@@ 5.1. Convergence test
Convergence of the centred scheme is tested using the relative $L^{2}$ norm between the initial and final solutions in the absence of stabilisation [Eldred2019]. Fig. 1(a) shows third and fourth order convergence for $p=1$ and $p=2$ finite elements respectively, for a range of spatial and temporal resolutions. The apparent super-convergence could be attributed to the fact that this test case is an extremely smooth test case, where no turbulent flow structures develop. For each line in Fig. 1, ${h}$ and ${\tau}$ reduce simultaneously at rate $C=\mathrm{CFL}/p^{2}$, where $\mathrm{CFL}=0.2$ for $p=1$ and $\mathrm{CFL}=0.1$ for $p=2$. Similar convergence rates are observed in Fig. 1(b) for the upwinded scheme.
@@ 5.2. Case study 1: thermal instability
We now consider transient case studies to confirm adequate conservation of system invariants under turbulent flow conditions and over long time periods relative to the onset of geostrophic turbulence. Discrete conservation is demonstrated using normalised values of discrete energy 15, entropy 18 and mass ($\mathcal{M}_{h}(t)=\int_{\Omega}\varphi_{h}$) on a logarithmic scale over time. The change in entropy due to forcing terms, $\partial\mathcal{S}_{h}/\partial t$ in 30, is also shown. The following numerical experiments correspond to $p=1$ finite elements and $\mathrm{CFL}=0.2$. Based on the convergence results evident in Fig. 1, using $p=2$ finite elements would require a smaller CFL value.
The analysis of conservation of system invariants in Sections 4.3 and 3.4 is based on a converged non-linear solution. We recognise other models solve with a fixed number of non-linear iterations per time step, as opposed to converging to a chosen tolerance [Maynard2020, Lee2024, Wood2013, Lee2024_helmholtz]. Such studies may also utilise preconditioning and multigrid methods to accelerate convergence [Maynard2020]. In the present study, the intent of the following numerical experiments is to demonstrate conservation of systems invariants in the presence of highly turbulent flow. So we accept non-linear solutions that converge to a tolerance of $10^{-12}$, meaning conservation errors up to $\mathcal{O}(10^{-12})$ are reasonable. A maximum of 50 non-linear iterations per time step is used in all simulations, even though the centred scheme generally converges in less than 20 non-linear iterations per time step.
The first transient case study involves a thermal instability [Eldred2019, Gouzien2017, Zeitlin2018, Kurganov2021]. Other studies simulate this test case using an upwinded finite volume scheme [Kurganov2021, Gouzien2017, Zeitlin2018]. A previous finite element study shows energy conservation at early times [Eldred2019], but entropy conservation, the evolution of complex non-linear flows, and the effect of upwinding is not considered. We extend these works using the novel centred and upwinded fluxes to simulate the thermal instability test case over a long time period while examining conservation of system invariants.
@@
Fig. 2 shows the evolution of the thermal instability for the centred and upwinded schemes. Similar to other studies that also consider this test case [Eldred2019, Gouzien2017], the initial growth evident in Fig. 2(a)–(d) reflects the wave number of 4 used in 38, and non-linear saturation of the instability occurs as the simulation progresses (Fig. 2(e)–(h)). Comparing the buoyancy field at $t=25$ (Fig. 2(a,b)) that obtained at $t=100$ (Fig. 2(g,h)) shows a large variation in buoyancy. Regardless, both the centred and upwinded schemes resolve the highly non-linear flow that develops as the simulation progresses. This is a direct consequence of the new linearised Jacobian 33, which allows for robust convergence for both centred and upwinded fluxes in the presence of significant variations from the initial mean flow state. As expected, including upwinded fluxes (Fig. 2(b,d,f,g)) suppresses spurious oscillations which arise in the absence of stabilisation (Fig. 2(a,c,e,g)). Therefore, we recommend using the upwinded scheme from the perspective of obtaining smooth solutions.
@@
We now assess the ability of the centred and upwinded schemes to conserve system invariants in the presence of turbulent flow. In the absence of stabilisation, Fig. 3(a,b) shows energy and mass conservation, while Fig. 3(c) demonstrates the entropy conservation error increases to $\mathcal{O}(10^{-8})$ over time even though ${\partial\mathcal{S}_{h}}/{\partial t}$ is machine zero for the entire duration (Fig. 3(d)). This supports the analysis in Section 4.3 above that entropy is not conserved pointwise in time. While an entropy conservation error of $\mathcal{O}(10^{-8})$ may seem large relative to energy and mass, it is comparable with other studies that consider temporal conservation of quadratic tracer invariants [Lee2022]. Thus, we conclude the centred scheme stably resolves turbulent dynamics over long periods of time and conserves system invariants.
For the upwinded scheme, the spikes in energy and mass evident in Fig. 3(a,b) correspond to instances when the non-linear solver fails to converge. In general, conservation errors obtained from numerical experiments are bounded by the tolerance of the non-linear solver. Consequently, non-linear solutions that fail to converge can lead to an increase in conservation error. As discussed above, we accept conservation errors up to $\mathcal{O}(10^{-12})$. The results in Fig. 3(a,b) demonstrate that failing to converge to the chosen non-linear tolerance leads to increased conservation errors of up to $\mathcal{O}(10^{-7})$ in energy and mass. While conservation errors of $\mathcal{O}(10^{-7})$ may seem large, such errors are comparable with other studies that choose to accept non-linear solutions after a fixed number of non-linear iterations per time step [Lee2024_helmholtz, Fig. 5].
@@
Values of $\epsilon=10^{-3}$ and $\epsilon=10^{-4}$ are chosen to approximate $\operatorname{sign}(x)$ in the upwinded scheme since $||\boldsymbol{F}_{h}^{n}\cdot\boldsymbol{n}^{+}||\approx 0.1$ for the centred scheme. When using the hard signum function with $\epsilon=10^{-4}$, solver convergence issues arise at $t\approx 45$. In contrast, using the soft signum function for $\epsilon=10^{-4}$ and $\epsilon=10^{-3}$ delays the onset of solver convergence issues to $t\approx 70$ and $t\approx 98$ respectively. Therefore, using the soft signum function and increasing $\epsilon$ assists solver convergence as turbulent flow dynamics evolve.
@@
Figure 3. Conservation errors for the thermal instability case study. (a,b,c) show normalised values of energy, mass and entropy, and (d) illustrates the change in entropy due to forcing terms 23. Each figure compares the centred and upwinded schemes for different approximations of signum and $\epsilon$ values. The vertical axis is logarithmic and the horizontal axis shows unit-less time. Parameters: $n_{s}=192$ spatial elements, $p=1$ spatial finite elements, $\mathrm{CFL}=0.2$.